# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ 2.3 The de Rham theorem for the $L^{2}$ de Rham complex
For $k=0,\dots,3$, the following relation holds:
@@
Since the classical version of de Rham’s theorem is formulated for the smooth complex, we give here a complete argument which uses a result of Costabel and McIntosh [14], suggested to us by Douglas Arnold.
@@
As an open subset of $\mathbb{R}^{3}$, $\Omega$ is a smooth manifold and the de Rham theorem applies, for instance as formulated in [24, Th. 18.14]. As a consequence, the $k$-chain cohomology groups (isomorphic to the dual spaces $(\mathcal{H}_{k})^{\prime}$ and hence of the same dimensions $\beta_{k}$) are isomorphic to the cohomology groups of the smooth de Rham complex
@@
Moreover, a Poincaré duality theorem holds ([24, Pbm. 18.7]), which states that the cohomology groups of the compactly supported de Rham sequence (indexed with reverse order) have the same dimensions. Namely, writing
@@ 2.4 Scalar and vector potentials
With regard to potentials, the question is then to know whether harmonic vector fields belong to some range space. It is clear from (19) that the tangential harmonic fields in $\mathfrak{H}^{1}$ have no scalar potential, from (20) we see that the normal harmonic fields in $\mathfrak{H}^{2}$ have no vector potential. But the reverse may be true, and a simple argument shows that it is indeed.
@@ 2.5 Scalar potentials for the normal harmonic fields
This problem is clearly well-posed thanks to Poincaré’s inequality on $H^{1}_{S}(\Omega)$, which holds since $S_{0}\neq\emptyset$, and we have the following result (see, e.g. [1, Prop. 3.18]).
@@
which is easy to verify: for all $1\leq j\leq\beta_{2}$, fix a function $\psi_{j}\in H^{1}_{S}(\Omega)$ which is constant on a (disconnected) neighborhood of $\partial\Omega$ and satisfies
@@ 3.1 Tunnel curves and reciprocal surfaces
Motivated by the discussion in [8, Sec. 3], our main assumption is twofold: that the domain admits piecewise smooth “tunnel curves” $\Gamma_{i}$, $1\leq i\leq\beta_{1}$, which form a basis for the 1-chain homology group $\mathcal{H}_{1}$, and that there exist surfaces $\Sigma_{i}$ which are reciprocal in a sense specified below. An important point is that these surfaces do not need to be disjoint.
Throughout the article we will say that a function $u$ defined on a bounded Lipschitz domain $M\subset\mathbb{R}^{3}$ is piecewise $C^{k}$ if there is a covering $\overline{M}=\cup_{a=1}^{m}\overline{M}_{a}$ into disjoint open cells $M_{a}$ with Lipschitz, piecewise $C^{1}$ boundaries, such that for all $1\leq a\leq m$, the restriction $u|_{M_{a}}$ is in $C^{k}(\overline{M_{a}})$. When useful, we will denote by $C^{k}(\mathcal{M})$ the space of such functions, $\mathcal{M}=(M_{a})_{a=1}^{m}$ being the associated covering of $M$.
@@
There exist closed curves $\Gamma_{i}\subset\partial\Omega$, and open oriented surfaces $\Sigma_{i}\subset\Omega$, $i=1,\ldots,\beta_{1}$, with the following properties:
  • for each $i$, there is a homeomorphism
@@
      the mapping $T_{i}$ is piecewise $C^{\infty}$ in the above sense.
  • each surface $\Sigma_{i}$ is an open part of a smooth manifold, with a possibly disconnected boundary $\partial\Sigma_{i}$ made of one or several closed curves $\tilde{\Gamma}_{i,n}$, $n=0,\dots,\beta_{1}(\Sigma_{i})$, which are piecewise smooth and included in $\partial\Omega$.
Finally, the curves and surfaces are reciprocal in the sense that $\Gamma_{i}$ intersects $\partial\Sigma_{j}$ once if $i=j$, and zero times if $i\neq j$. More precisely, observing that $\Gamma_{i}\cap\partial\Sigma_{j}\subset X_{i}\cap\partial\Sigma_{j}$, we assume that
$X_{i}\cap\partial\Sigma_{j}=\begin{cases}T_{j}(\{0\}\times\{0\}\times(-1,1))&\text{ if }i=j,\\ \emptyset&\text{ if }i\neq j.\end{cases}$ (31)
$\mathcal{O}=[0,1)\times\mathbb{T}\times(-1,1)$
[…]
$\Omega$
$T_{i}:\mathcal{O}\to X_{i}$
[…]
Figure 2: Parametrization of a tunnel curve $\Gamma_{i}$ and its neighborhood $X_{j}$ using a piecewise smooth homeomorphism $T_{j}$ defined on a (straight) solid torus $\mathcal{O}$ periodic along $\theta$.
@@
The surfaces $\Sigma_{i}$ are reminiscent of the cut surfaces used to construct piecewise smooth scalar potentials in [18, 31, 1], and in the analysis they also allows us to show the linear independence of the resulting harmonic fields. One difference is that our construction does not require them to be disjoint. We note that it does not seem possible to find disjoint cut surfaces for the domain shown in Figure 1, and more generally for a hollow torus.
@@ 3.2 The surface flux functionals
Our assumptions on the surfaces $\Sigma_{i}$ guarantee that they are surface with corners [24], and more generally standard surfaces in the sense of [36, Sec. III-18], so that the Stokes theorem applies. Using the correspondence between vector fields and 1-forms, this yields
$\int_{\Sigma_{i}}{\bm{n}}\cdot\operatorname{{\bf curl}}{\bm{u}}=\int_{\partial\Sigma_{i}}{\bm{\tau}}\cdot{\bm{u}}$ (32)
for any vector field ${\bm{u}}\in C^{1}(\overline{\Omega})^{3}$. We observe that (32) also holds for ${\bm{u}}\in H(\operatorname{{\bf curl}};\Omega)$ which is piecewise $C^{1}$ in the sense specified above. Indeed, for every interface $f=\overline{M_{a}}\cap\overline{M_{b}}$ between two subdomains, and every vector ${\bm{\tau}}_{f}$ tangent to $f$, such a field satisfies
${\bm{\tau}}_{f}\cdot{\bm{u}}|_{M_{a}}={\bm{\tau}}_{f}\cdot{\bm{u}}|_{M_{b}}\quad\text{ on }f.$ (33)
Decomposing $\Sigma_{i}$ into subsurfaces $\Sigma_{i,a}=\Sigma_{i}\cap M_{a}$ with piecewise $C^{1}$ boundaries, one then applies the Stokes formula on each of these subsurfaces which yields a collection of subsurface integrals which sum to the surface integral in (32), and two types of line integrals: line integrals corresponding to interior curves (shared by two subsurfaces) $\gamma_{i,a,b}=\partial\Sigma_{i,a}\cap\partial\Sigma_{i,b}$ and line integrals over boundary curves $\gamma_{i,a}=\partial\Sigma_{i,a}\cap\partial\Sigma$. Since the former ones appear twice with opposite orientations, their sum vanishes thanks to (33), and the latter ones sum to the boundary integral in (32).
Using the same arguments as in the proof of [1, Lemma 3.10], which rely on the properties of $\Sigma_{i}$ listed in Assumption 3.1, we can extend the mapping $\gamma_{n}:C^{0}_{c}(\Omega)\to L^{\infty}(\Sigma_{i}),{\bm{w}}\mapsto{\bm{n}}\cdot{\bm{w}}|_{\Sigma_{i}}$ to a bounded trace operator $\gamma_{n}:H_{0}(\operatorname{div};\Omega)\to(H^{1/2}(\Sigma_{i}))^{\prime}.$ For each $1\leq i\leq\beta_{1}$, this allows us to consider a flux functional on $\Sigma_{i}$ defined as
$\Phi_{i}:H_{0}(\operatorname{div};\Omega)\to\mathbb{R},\quad{\bm{w}}\mapsto\int_{\Sigma_{i}}{\bm{n}}\cdot{\bm{w}}.$ (34)
@@
The linear form $\Phi_{i}$ vanishes on $\operatorname{{\bf curl}}H_{0}(\operatorname{{\bf curl}};\Omega)$, i.e., we have
$\Phi_{i}(\operatorname{{\bf curl}}{\bm{v}})=0,\quad{\bm{v}}\in H_{0}(\operatorname{{\bf curl}};\Omega).$ (35)
@@
For ${\bm{v}}\in C^{\infty}_{c}(\Omega)$, Stokes’ formula (32) yields
$\Phi_{i}(\operatorname{{\bf curl}}{\bm{v}}_{n})=\int_{\Sigma_{i}}{\bm{n}}\cdot\operatorname{{\bf curl}}{\bm{v}}_{n}=\int_{\partial\Sigma_{i}}{\bm{\tau}}\cdot{\bm{v}}_{n}=0$ (36)
where the last equality follows from $\partial\Sigma_{i}\subset\partial\Omega$. We conclude with a density argument: given a sequence ${\bm{v}}_{n}\in C^{\infty}_{c}(\Omega)$ converging to an arbitrary $\bm{v}$} in $H_{0}(\operatorname{{\bf curl}};\Omega)$, we have $\operatorname{{\bf curl}}{\bm{v}}_{n}\to\operatorname{{\bf curl}}{\bm{v}}$ in $H_{0}(\operatorname{div};\Omega)$ and the result follows from the continuity of the operator (34). ∎

### v2

@@ 2.3 The de Rham theorem for the $L^{2}$ de Rham complex
For $k=0,\dots 3$, the following relation holds:
@@
Since the classical version of de Rham’s theorem is formulated for the smooth complex, we give here a detailed argument which uses a result of Costabel and McIntosh [18], suggested to us by Douglas Arnold.
@@
According to Remark 2.1 we may consider the homology groups on $\Omega$ which, as an open subset of $\mathbb{R}^{3}$, is a smooth manifold (without boundary). We can then apply the de Rham theorem as formulated in [30, Th. 18.14], which states that the $k$-chain cohomology groups $\mathcal{H}^{k}(\Omega;\mathbb{R})$ (which are real vector spaces isomorphic to the dual spaces $(\mathcal{H}_{k}(\Omega))^{\prime}$ of linear forms on $\mathcal{H}_{k}(\Omega)$, and hence of the same dimensions $\beta_{k}$) are isomorphic to the cohomology groups (spaces) of the smooth de Rham complex
@@
Moreover, a Poincaré duality theorem holds ([30, Pbm. 18.7]), which states that the cohomology groups of the compactly supported de Rham sequence (indexed with reverse order) have the same dimensions. Namely, writing
@@ 2.4 Scalar and vector potentials
With regard to potentials, the question is then to know whether harmonic vector fields belong to some range space. It is clear from (19) that the tangential harmonic fields in $\mathfrak{H}^{1}$ have no scalar potential, and from (20) we see that the normal harmonic fields in $\mathfrak{H}^{2}$ have no vector potential. But the reverse may be true, and a simple argument shows that it is indeed.
@@
Remark 2.3.
For a characterization of smooth fields with scalar or vector potentials, see [16].
@@ 2.5 Scalar potentials for the normal harmonic fields
This problem is clearly well-posed thanks to Poincaré’s inequality on $H^{1}_{S}(\Omega)$, which holds since $S_{0}\neq\emptyset$, and we have the following result (see, e.g. [3, Prop. 3.18]).
@@
which is easy to verify: for all $1\leq j\leq\beta_{2}$, fix a function $\psi_{j}\in H^{1}_{S}(\Omega)$ that is constant on a (disconnected) neighborhood of $\partial\Omega$ and satisfies
@@ 3.1 Tunnel curves and reciprocal surfaces
Motivated by the discussion in [11, Sec. 3], our construction relies on the assumption that the domain $\Omega$ admits piecewise smooth “tunnel curves” $\Gamma_{i}$, which are closed curves lying on the boundary $\partial\Omega$, and “reciprocal surfaces” $\Sigma_{i}\subset\overline{\Omega}$, $1\leq i\leq\beta_{1}$, whose boundaries $\partial\Sigma_{i}$ also lie on the domain boundary and intersect the tunnel curves in a dual manner, in the sense that the number of intersection points of $\Gamma_{i}$ and $\Sigma_{j}$ satisfies
$I(\Gamma_{i},\Sigma_{j})=\delta_{i,j},\quad 1\leq i,j\leq\beta_{1}.$ (28)
Aside from regularity considerations, the fact that such a property holds for general domains follows from classical results in algebraic topology, and in particular from the geometric realization of Poincaré-Lefschetz duality through intersection pairings. This important duality result states that the $k$-chain homology groups $\mathcal{H}_{k}(\overline{\Omega})$ are isomorphic to the (3k)(3-k)-chain relative homology groups $\mathcal{H}_{3-k}(\overline{\Omega},\partial\Omega)$, which are the homology groups of the singular chains modulo the boundary – two chains being equal modulo the boundary if their difference lies on $\partial\Omega$, see e.g. [31, 11]. A modern exposition of these results can be found in the detailed lecture notes of Stefan Friedl [23]: applying Theorem 194.14 and Proposition 198.2 one finds indeed that there exists a non-singular pairing between the above (torsion-free) homology groups,
$Q=Q^{as}_{\overline{\Omega},\emptyset,\partial\Omega}:\mathcal{H}_{k}(\overline{\Omega})\times\mathcal{H}_{3-k}(\overline{\Omega},\partial\Omega)\to\mathbb{Z}~,$ (29)
called asymmetric intersection pairing for the reason that for any pair of cycles $\gamma\in\mathcal{Z}_{k}(\overline{\Omega})$ and $\sigma\in\mathcal{Z}_{3-k}(\overline{\Omega},\partial\Omega)$ which intersect nicely, $Q([\gamma],[\sigma])$ coincides with the algebraic intersection number of $\gamma$ and $\sigma$.
With $k=1$, we thus find that for any collection of (reasonably smooth) closed curves $\Gamma_{i}$ representing a basis of $\mathcal{H}_{1}(\overline{\Omega})$, there exists a collection of surfaces $\Sigma_{i}$ which are closed modulo to the boundary, represent a basis of $\mathcal{H}_{2}(\overline{\Omega},\partial\Omega)$ and are such that $Q([\Gamma_{i}],[\Sigma_{j}])=\delta_{i,j}$ for all $i,j$. We then observe that the curves $\Gamma_{i}$ may be chosen to lie on the domain boundary without changing their homology class $\Gamma_{i}]$, and that the surfaces fulfill the above requirements (as long as they can be chosen to intersect nicely the curves, which is not a restrictive assumption in a Lipschitz domain $\Omega$): being closed modulo the boundary means indeed that $\partial\Sigma_{i}\subset\partial\Omega$, and a relation of the form (28) follows directly from the $Q$-duality of the homology classes, with the specification that $I$ denotes the algebraic intersection number of the singular chains, which weighs each intersection with a sign corresponding to their respective orientations.
$\mathcal{O}=[0,1)\times\mathbb{T}\times(-1,1)$
[…]
$\Omega$
$T_{i}:\mathcal{O}\to X_{i}$
[…]
Figure 2: Illustration of a tunnel curve $\Gamma_{i}$ and its neighborhood $X_{i}$, parametrized with a piecewise smooth homeomorphism $T_{i}$ defined on the (straight) torus $\mathcal{O}$, periodic along the variable $\theta$.
To formulate our precise working hypotheses, we now introduce a couple of definitions.
Definition 3.1.
A function $u$ defined on a bounded Lipschitz domain $M\subset\mathbb{R}^{d}$ is said piecewise $C^{k}$, $k\geq 0$, if there exists a decomposition $\overline{M}=\cup_{a=1}^{m}\overline{M}_{a}$ into disjoint open cells $M_{a}$ with Lipschitz, piecewise $C^{1}$ boundaries, such that for all $1\leq a\leq m$, the restriction $u|_{M_{a}}$ is in $C^{k}(\overline{M_{a}})$. If in addition $u$ is invertible and $u^{-1}$ is piecewise $C^{k}$ over the cells $u(M_{a})$, $1\leq a\leq m$, then we say that $u$ is a piecewise $C^{k}$ diffeomorphism.
In the sequel we will refer to such a decomposition $\mathcal{M}=\{M_{a}:a=1,\dots m\}$ as a mesh, and we will denote by $C^{k}(\mathcal{M})$ the space of piecewise $C^{k}$ functions over $\mathcal{M}$.
Definition 3.2.
We say that a surface $\Sigma$ and a mesh $\mathcal{M}$ are compatible if there exists a decomposition $\Sigma=\cup_{a=1}^{n}\Sigma_{a}$ made of smooth surfaces with corners [30], such that the interiors $\Sigma_{a}\setminus\partial\Sigma_{a}$ are disjoint and included in closed mesh cells $\overline{M}_{b}$ for some $b=b(a)\in\{0,\dots m\}$.
Our working hypotheses read then as follows.
@@
There exist oriented curves $\Gamma_{i}\subset\partial\Omega$ and surfaces $\Sigma_{i}\subset\overline{\Omega}$, $i=1,\ldots\beta_{1}$, with the following properties:
  • Each curve $\Gamma_{i}$ is closed and piecewise smooth, in particular there is a homeomorphism
@@
      and $\Gamma_{i}$ is oriented by the unit tangent vector ${\bm{\tau}}_{\Gamma_{i}}$ along increasing values of $\theta$
@@
      the mapping $T_{i}$ is a piecewise $C^{\infty}$ diffeomorphism in the sense of Definition 3.1. Letting $\mathcal{O}_{i,a}$, $a=1,\dots m_{i}$, be the associated open mesh cells of $\mathcal{O}$, we denote by
      $\mathcal{X}_{i}=\{X_{i,a}:a=0,\dots m_{i}\}\quad\text{with}\quad\begin{cases}X_{i,0}:=\Omega\setminus\overline{X}_{i},\\ X_{i,a}:=T_{i}(\mathcal{O}_{i,a}),~~a=1,\dots m_{i},\end{cases}$ (33)
      the corresponding mesh of $\Omega$.
  • Each surface $\Sigma_{i}$ is Lipschitz, compatible with all the meshes $\mathcal{X}_{j}$, $j=1,\dots\beta_{1}$, in the sense of Definition 3.2, and its (manifold) boundary $\partial\Sigma_{i}$ is included in $\partial\Omega$.
Finally, the above curves and surfaces are reciprocal in the sense that $\Gamma_{i}$ intersects $\Sigma_{j}$ once if $i=j$, and zero times if $i\neq j$. More precisely, observing that $\Gamma_{i}\cap\Sigma_{j}=\Gamma_{i}\cap\partial\Sigma_{j}\subset X_{i}\cap\partial\Sigma_{j}$, it holds
$X_{i}\cap\partial\Sigma_{j}=\begin{cases}T_{i}\big(\{0\}\times\{0\}\times(-1,1)\big)&\text{ if }i=j,\\ \emptyset&\text{ if }i\neq j,\end{cases}$ (34)
and we orient each surface $\Sigma_{i}$ by letting its unit normal vector ${\bm{n}}_{\Sigma_{i}}$ at the intersection point $P_{i}=\Gamma_{i}\cap\Sigma_{i}=T_{i}(0,0,0)$ satisfy ${\bm{n}}_{\Sigma_{i}}\cdot{\bm{\tau}}_{\Gamma_{i}}>0$.
@@
In the above assumption, and in particular in (34), we have considered for simplicity that an intersection duality of the form (28) can be realized with geometric intersection numbers $I(\Gamma_{i},\Sigma_{j})=\#(\Gamma_{i}\cap\Sigma_{j})$, which is the case for virtually any domain we can think of. In the general case however, this duality is only established with algebraic intersection numbers as discussed above. Instead of (34), one should thus assume that for any $i,j$, there exists distinct points $\theta(i,j,p)\in\mathbb{T}$, $1\leq p\leq n(i,j)$ (with $n(i,j)$ possibly zero), such that
$X_{i}\cap\partial\Sigma_{j}=\cup_{1\leq p\leq n(i,j)}\,T_{i}\big(\{0\}\times\{\theta(i,j,p)\}\times(-1,1)\big)$ (35)
holds, with normal vectors ${\bm{n}}_{\Sigma_{i}}$ and indices at the intersection points $P_{i,j,p}=T_{i}(0,\theta(i,j,p),0)$,
${\rm ind}(i,j,p):={\rm sign}\Big(({\bm{\tau}}_{\Gamma_{i}}\cdot{\bm{n}}_{\Sigma_{j}})(P_{i,j,p})\Big)\in\{-1,+1\},$ (36)
that satisfy the duality relation
$I(\Gamma_{i},\Sigma_{j}):=\sum_{p=1}^{n(i,j)}{\rm ind}(i,j,p)=\delta_{i,j},\quad 1\leq i,j\leq\beta_{1}.$ (37)
In our analysis below we will consider this general case since it is handled by the same arguments.
Remark 3.5.
The fact that the curves $\Gamma_{i}$ represent a basis of the first homology group $\mathcal{H}_{1}(\overline{\Omega})$ follows from Assumption 3.3 and properties of the intersection pairing (29) recalled at the beginning of Section 3.1. Indeed, these curves being 1-cycles they represent 1-chain homology classes, and similarly the surfaces $\Sigma_{i}$ being closed modulo the boundary they represent relative homology classes in $\mathcal{H}_{2}(\overline{\Omega},\partial\Omega)$. Their intersection duality assumed in (34) (or more generally in (35)–(37)) then yields $Q([\Gamma_{i}],[\Sigma_{j}])=\delta_{i,j}$, which shows the linear independence of the classes $\Gamma_{i}]$. Their generating property follows from the definition of $\beta_{1}$.
@@ 3.2 The surface flux functionals
The main purpose of the reciprocal surfaces $\Sigma_{i}$ is to establish the linear independence of our tangent harmonic vector fields. To this end we first review a version of the usual Stokes formula for piecewise smooth curl-conforming fields.
@@
Let $\Sigma$ be an oriented surface compatible with a mesh $\mathcal{M}$ in the sense of Definition 3.2. Then the Stokes formula
$\int_{\Sigma}{\bm{n}}\cdot\operatorname{{\bf curl}}{\bm{u}}=\int_{\partial\Sigma}{\bm{\tau}}\cdot{\bm{u}}$ (38)
holds for any ${\bm{u}}\in C^{\infty}(\mathcal{M})\cap H(\operatorname{{\bf curl}};\Omega)$.
@@
Since any surface $\Sigma_{a}$, $a=1,\dots n$, involved in the decomposition of Definition 3.2 is a smooth surface with corners and is included in some domain $\overline{M}_{b(a)}$ where $\bm{u}$} is smooth, Stokes’ theorem applies. Using the correspondence between vector fields and differential forms (see for instance [30, Ch. 16]), this yields
$\int_{\Sigma_{a}}{\bm{n}}\cdot\operatorname{{\bf curl}}{\bm{u}}=\int_{\partial\Sigma_{a}}{\bm{\tau}}\cdot{\bm{u}}|_{M_{b(a)}},\qquad 1\leq a\leq n.$ (39)
Summing over $a$ then yields a collection of surface integrals which sum to the left-hand side integral in (38), and a collection of line integrals on the right-hand side: some corresponding to interior curves $\gamma_{a,a^{\prime}}=\partial\Sigma_{a}\cap\partial\Sigma_{a^{\prime}}$ with $a\neq a^{\prime}$, and others corresponding to boundary curves $\gamma_{a}=\partial\Sigma_{a}\cap\partial\Sigma$. On interior curves which appear twice with opposite orientations, we have
$\int_{\gamma_{a,a^{\prime}}}{\bm{\tau}}_{\partial\Sigma_{a}}\cdot{\bm{u}}|_{M_{b(a)}}+\int_{\gamma_{a,a^{\prime}}}{\bm{\tau}}_{\partial\Sigma_{a^{\prime}}}\cdot{\bm{u}}|_{M_{b(a^{\prime})}}=\int_{\gamma_{a,a^{\prime}}}{\bm{\tau}}_{\partial\Sigma_{a}}\cdot({\bm{u}}|_{M_{b(a)}}-{\bm{u}}|_{M_{b(a^{\prime})}})=0$
where the last equality follows from the continuity of the tangential traces of piecewise smooth, globally $H(\operatorname{{\bf curl}};\Omega)$ fields, see e.g. [8, Prop. 2.2.32]. As a result, only the line integrals over boundary curves remain in the sum, yielding the right-hand side integral in (38). ∎
Let us further study the flux functionals defined on the surfaces $\Sigma_{i}$.
Proposition 3.7.
For each $1\leq i\leq\beta_{1}$, the flux functional
$\Phi_{i}:C^{\infty}_{c}(\Omega)\to\mathbb{R},\qquad{\bm{w}}\mapsto\int_{\Sigma_{i}}{\bm{n}}\cdot{\bm{w}}.$ (40)
extends to a continuous linear form on $H_{0}(\operatorname{div};\Omega)$ which vanishes on $\operatorname{{\bf curl}}H_{0}(\operatorname{{\bf curl}};\Omega)$: we have
$\Phi_{i}(\operatorname{{\bf curl}}{\bm{u}})=0,\qquad{\bm{u}}\in H_{0}(\operatorname{{\bf curl}};\Omega).$ (41)
Proof.
The argument is similar to that used in the proof of [3, Lemma 3.10], although simpler since we consider fluxes defined on single surfaces. Using the fact that $\Sigma_{i}$ is Lipschitz and oriented, we can find a scalar function ${\varphi}_{i}\in H^{1}(\omega_{i})$ with $\omega_{i}=\Omega\setminus\Sigma_{i}$, whose trace vanishes on the positive side of $\Sigma_{i}$ (as pointed by its normal vector $\bm{n}$}) and is 1 on its negative side. For ${\bm{w}}\in C^{\infty}_{c}(\Omega)$, Green’s formula yields then
$\Phi_{i}({\bm{w}})=\int_{\Sigma_{i}}{\bm{n}}\cdot{\bm{w}}=\int_{\omega_{i}}{\bm{w}}\cdot\operatorname{{\bf grad}}\phi_{i}+\int_{\omega_{i}}(\operatorname{div}{\bm{w}})\phi_{i}$ (42)
and in particular one has
$\Phi_{i}({\bm{w}})\leq\lVert{\bm{w}}\rVert_{L^{2}(\omega_{i})}\lVert\operatorname{{\bf grad}}\phi_{i}\rVert_{L^{2}(\omega_{i})}+\lVert\operatorname{div}{\bm{w}}\rVert_{L^{2}(\omega_{i})}\lVert\phi_{i}\rVert_{L^{2}(\omega_{i})}\leq\lVert{\bm{w}}\rVert_{H(\operatorname{div};\Omega)}\lVert\phi_{i}\rVert_{H^{1}(\omega_{i})}$ (43)
so that $\Phi_{i}$ admits indeed a continous extension in $H_{0}(\operatorname{div};\Omega)$. To verify (41) we then observe that for ${\bm{u}}\in C^{\infty}_{c}(\Omega)$, Stokes’ formula (38) gives
$\Phi_{i}(\operatorname{{\bf curl}}{\bm{u}})=\int_{\Sigma_{i}}{\bm{n}}\cdot\operatorname{{\bf curl}}{\bm{u}}=\int_{\partial\Sigma_{i}}{\bm{\tau}}\cdot{\bm{u}}=0$ (44)
where the last equality follows from the inclusion $\partial\Sigma_{i}\subset\partial\Omega$. We conclude by using the density of $C^{\infty}_{c}(\Omega)$ in $H_{0}(\operatorname{{\bf curl}};\Omega)$, the inclusion $\operatorname{{\bf curl}}H_{0}(\operatorname{{\bf curl}};\Omega)\subset H_{0}(\operatorname{div};\Omega)$, and the continuity of $\Phi_{i}$ in the latter space. ∎